Bed & Bath, a retailing company, has two departments—Hardware and Linens. The company’s most recent monthly contribution format income statement follows:

Department
Total Hardware Linens
Sales $ 4,100,000 $ 3,080,000 $ 1,020,000
Variable expenses 1,292,000 882,000 410,000
Contribution margin 2,808,000 2,198,000 610,000
Fixed expenses 2,150,000 1,310,000 840,000
Net operating income (loss) $ 658,000 $ 888,000 $ (230,000 )

A study indicates that $372,000 of the fixed expenses being charged to Linens are sunk costs or allocated costs that will continue even if the Linens Department is dropped. In addition, the elimination of the Linens Department will result in a 13% decrease in the sales of the Hardware Department.

Required:

What is the financial advantage (disadvantage) of discontinuing the Linens Department?
Answer #1

Contribution margin for Hardware would be=2,198,000*(1-Decrease in Growth)

=2,198,000*(1-0.13)=1912260

Less:Fixed cost for Hardware=(1.310,000)

Net operating income for Hardware=$602260

Less:Unavoidable fixed cost of Linens=(372,000)

Total net operating income=$230260

Hence financial disadvantage=(658,000-230260)

=$427740